Use the slope formula to find the value of x that passes through the given point.
(x,7);(8,11), m=4/3

Answer by lwsshak3(11628)

Use the slope formula to find the value of x that passes through the given point.
(x,7);(8,11), m=4/3
***
equation of a line: y=mx+b, m=slope, b=y-intercept
For given problem:
y=4x/3+b
solve for b using coordinates of given point(8,11):
11=4*8/3+b
b=33/3-32/3=1/3
b=1/3
equation: y=4x/3+1/3
solving for x on given line:
7=4x/3+1/3
lcd:3
21=4x+1
4x=20
x=5
value of x that passes through the given point=5